determine the critical points of f(x)=8x3+81x2-42x-8


1
Expert's answer
2021-12-23T16:02:18-0500
"f(x)=8x^3+81x^2-42x-8"

Domain: "(-\infin, \infin)"

Find the first derivative


"f'(x)=24x^2+162x-42"

Find the critical number(s)


"f'(x)=0=>24x^2+162x-42=0"

"4x^2+27x-7=0"

"D=(27)^2-4(4)(-7)=841"

"x=\dfrac{-27\pm\sqrt{841}}{2(4)}"

"x_1=\dfrac{-27-29}{8}=-7"

"x_2=\dfrac{-27+29}{8}=\dfrac{1}{4}"

Critical numbers: "-7, \dfrac{1}{4}."


If "x<-7, f'(x)>0, f(x)" increases.

If "-7<x<\dfrac{1}{4}, f'(x)<0, f(x)" decreases.

If "x>\dfrac{1}{4}, f'(x)>0, f(x)" increases.


"f(-7)=8(-7)^3+81(-7)^2-42(-7)-8=1511"

"f(\dfrac{1}{4})=8(\dfrac{1}{4})^3+81(\dfrac{1}{4})^2-42(\dfrac{1}{4})-8=-\dfrac{213}{16}"

Local maximum: "(-7, 1511)"


Local minimum: "(\dfrac{1}{4}, -\dfrac{213}{16})"